OpenStudy (anonymous):

Douglas has a credit card with an interest rate of 11.05%, compounded monthly. He used his credit card to buy a new sofa, which cost $670 before the sales tax of 7.94%. Douglas paid off his balance by making equal monthly payments for three years. Assuming that he had no other purchases on his credit card, how much did Douglas pay in total for the sofa? (Round all dollar values to the nearest cent.)

OpenStudy (nethercreep333):

The amount financed is 670 x (1+0.0794) = $723.20 Next, find the monthly payments ... P[ 1 - (1+0.1105/12)^(-12*3)] / (0.1105/12) = 723.20 P = $23.69/mo Total payments = 23.69 x 36 = $852.84

OpenStudy (nethercreep333):

medal me